a) Determine the probability distribution for the number of women on a civil- court jury of 6, selected from a pool of 8 men and 10 women
b) what is the probability that there are at least 2 women in the jury ?
1
Expert's answer
2020-03-16T12:10:37-0400
"\binom{18}{6}={18! \over 6!(18-6)!}=18564"


"\binom{8}{0}={8! \over 0!(18-0)!}=1"

"\binom{8}{1}={8! \over 1!(18-1)!}=8"

"\binom{8}{2}={8! \over 2!(18-2)!}=28"


"\binom{8}{3}={8! \over 3!(18-3)!}=56"

"\binom{8}{4}={8! \over 4!(18-4)!}=70"

"\binom{8}{5}={8! \over 5!(18-5)!}=56"

"\binom{8}{6}={8! \over 6!(18-6)!}=28"

"\binom{10}{0}={10! \over 0!(10-0)!}=1"

"\binom{10}{1}={10! \over 1!(10-1)!}=10"

"\binom{10}{2}={10! \over 2!(10-2)!}=45"

"\binom{10}{3}={10! \over 3!(10-3)!}=120"

"\binom{10}{4}={10! \over 4!(10-4)!}=210"

"\binom{10}{5}={10! \over 5!(10-5)!}=252"

"\binom{10}{6}={10! \over 6!(10-6)!}=210"

Let "X=" the number of women on a civil- court jury of 6.


"P(X=x)={\dbinom{10}{x}\dbinom{8}{6-x} \over \dbinom{18}{6}}"

"P(X=0)={\dbinom{10}{0}\dbinom{8}{6-0} \over \dbinom{18}{6}}={1(28) \over 18564}={1 \over 663}"

"P(X=1)={\dbinom{10}{1}\dbinom{8}{6-1} \over \dbinom{18}{6}}={10(56) \over 18564}={20 \over 663}"

"P(X=2)={\dbinom{10}{2}\dbinom{8}{6-2} \over \dbinom{18}{6}}={45(70) \over 18564}={75 \over 442}"

"P(X=3)={\dbinom{10}{3}\dbinom{8}{6-3} \over \dbinom{18}{6}}={120(56) \over 18564}={80 \over 221}"

"P(X=4)={\dbinom{10}{4}\dbinom{8}{6-4} \over \dbinom{18}{6}}={210(28) \over 18564}={70 \over 221}"

"P(X=5)={\dbinom{10}{5}\dbinom{8}{6-5} \over \dbinom{18}{6}}={252(8) \over 18564}={24 \over 221}"

"P(X=6)={\dbinom{10}{6}\dbinom{8}{6-6} \over \dbinom{18}{6}}={210(1) \over 18564}={5 \over 442}"

Check


"{1 \over 663}+{20 \over 663}+{75 \over 442}+{80 \over 221}+{70 \over 221}+{24 \over221}+{5 \over 442}=1"

The probability distribution for the number of women on a civil- court jury of 6


"\def\arraystretch{1.5}
 \begin{array}{c: c: c: c: c: c:c: c: c: c}
 x_k & 0 & 1 & 2 & 3 & 4 & 5 & 6 \\ \hline
 f(x_k) & {1 \over 663} & {20 \over 663} & {75 \over 442} & {80\over 221} & {70 \over 221} & {24 \over 221} & {5 \over 442} 
\end{array}"

b) The probability that there are at least 2 women in the jury is


"P(X\geq2)=1-P(X=0)-P(X=1)="

"=1-{1 \over 663}-{20 \over 663}={214 \over 221}"